# A Low-Cost, Low-Power Media Converter Solution for Next-Generation Detector Readout Systems

### v3

@@ 6 Achievements and Future Directions
The proposed solution’s cost is compared to the high-end FPGA readout board used in LHCb, the PCIe400, which processes 46 lpGBT links and is priced at \$16,600 [12], resulting in a cost of \$360 per lpGBT link. Most of the PCIe400 cost is attributed to the FPGA, with an estimated \$10,000 for the FPGA and \$6,600 for the rest of the board. Assuming similar costs for optical transceivers222four 12-link Samtec FireFly modules [13] leaves \$4,200 for manufacturing and remaining bill of material costs.
The next iteration of the proposed solution will handle 48 lpGBT links. The selected SoM is the Enclustra XRU90, priced at \$4,123[14]. The SoM absorbs most board costs, and hosting multiple SoMs on the same baseboard enables sharing of expensive components like low-jitter PLLs and voltage regulators. Based on these considerations and the cost of existing SoM baseboards, we estimate an additional \$500 per SoM for baseboard costs.
This analysis demonstrates that the proposed solution is highly cost-effective, with a cost of \$134 per lpGBT link compared to existing designs, as summarized in Table 2.
Board PCIe400 NetGBT
FPGA SoM Cost 10000 4123
Transceivers Cost 2400 2400
PCB Cost 4200 500
Total Board Cost 16600 6423
lpGBT Links 46 48
Cost per lpGBT link 361 134
Table 2: Estimated cost comparison between the PCIe400 FPGA readout board and the NetGBT solution. All prices are in \$USD.

### v5

@@ 6 Achievements and Future Directions
The next iteration of the proposed solution will support 48 lpGBT links. The selected System-on-Module (SoM) is the Enclustra XRU90, priced at \$4,123 [12] for a single unit, with potential cost reductions at higher volumes. For optical transceivers, a 12-link Samtec FireFly™ module is listed at \$920 from the distributor [13], resulting in a cost of \$77 per link. By comparison, a COTS Ethernet transceiver is priced at \$40 [14], resulting in \$10 per link, suggesting the realistic cost for optical links can be around \$43 per link. Using receive-only transceiver modules could further reduce costs to about one-third of this value. Additionally, 100GbE transceivers for the uplink are priced at \$43 [15].
The SoM integrates most of the board’s cost, and using multiple SoMs on a shared baseboard allows for the distribution of expensive components, such as low-jitter PLLs and voltage regulators. Based on these factors and the pricing of existing SoM baseboards, an additional \$500 per SoM is estimated for baseboard costs.
This evaluation highlights the cost-effectiveness of the proposed solution, achieving a total cost of \$142 per lpGBT link (Table 2).
FPGA SoM Cost 4100
lpGBT Transceivers Cost 2000
100GbE Transceivers Cost 200
PCB Cost 500
Total Board Cost 6800
Cost per lpGBT link 142
Table 2: Estimated cost of the NetGBT solution capable of handling 48 lpGBT links. All prices are in \$USD.